# HumorDB: Can AI understand graphical humor?

### v2

@@ 3.2 Assessing human performance
    Outlier Removal: Participants with $>$10% of ratings having z-scores outside $\pm$1.96 were excluded
@@
  • Binary and Comparison Tasks: Excluded participants with $>$3 inconsistent responses on repeated images
  • Range Task: Excluded participants if any single image received ratings differing by $>$4 points, or if $>$4 images showed differences exceeding 2 points
@@
[…]
@@
[…]
@@
[…]
@@
Training 2,1362,1362 , 136 698 1,068 0.79±0.180.79$\pm$ 0.180.79 ± 0.18 5.75±1.425.75$\pm$ 1.425.75 ± 1.42 1,068 0.12±0.160.12$\pm$ 0.160.12 ± 0.16 3.62±1.753.62$\pm$ 1.753.62 ± 1.75
Valid. 703703703 273 351 0.78±0.190.78$\pm$ 0.190.78 ± 0.19 5.68±1.405.68$\pm$ 1.405.68 ± 1.40 351 0.12±0.170.12$\pm$ 0.170.12 ± 0.17 3.65±1.653.65$\pm$ 1.653.65 ± 1.65
testAll 706706706 300 352 0.77±0.190.77$\pm$ 0.190.77 ± 0.19 5.60±1.375.60$\pm$ 1.375.60 ± 1.37 352 0.14±0.170.14$\pm$ 0.170.14 ± 0.17 3.39±1.683.39$\pm$ 1.683.39 ± 1.68
testOnlyPairs 600600600 300 300 0.77±0.190.77$\pm$ 0.190.77 ± 0.19 5.62±1.365.62$\pm$ 1.365.62 ± 1.36 300 0.13±0.160.13$\pm$ 0.160.13 ± 0.16 3.30±1.683.30$\pm$ 1.683.30 ± 1.68
Total 3,542 1,271 1,771 0.79±0.200.79$\pm$ 0.200.79 ± 0.20 5.70±1.405.70$\pm$ 1.405.70 ± 1.40 1,771 0.13±0.170.13$\pm$ 0.170.13 ± 0.17 3.58±1.723.58$\pm$ 1.723.58 ± 1.72
@@ 5 Results
Explanation Quality. Explanation Quality. Our human evaluation of VLM explanations revealed strong performance, particularly from larger models. On a 1–5 scale, Gemini’s explanations were rated highest for both identifying the source of humor (4.1±0.64.1$\pm$ 0.64.1 ± 0.6) and for overall accuracy (4.24±0.474.24$\pm$ 0.474.24 ± 0.47). GPT-4o also performed well (Why Funny: 3.9±0.73.9$\pm$ 0.73.9 ± 0.7, Accuracy: 4.12±0.634.12$\pm$ 0.634.12 ± 0.63), followed by LLaVA (Why Funny: 3.5±0.93.5$\pm$ 0.93.5 ± 0.9, Accuracy: 3.57±0.953.57$\pm$ 0.953.57 ± 0.95). These results suggest a robust, though not perfect, ability to articulate visual humor beyond simple keyword matching. Automated scores are detailed in Appendix 8.
@@
Model Name testAllSet RMSE ±$\pm SD$ testOnlyPairs RMSE ±$\pm SD$
dinov2 large 1.98±0.081.98$\pm$ 0.081.98 ± 0.08 1.96±0.091.96$\pm$ 0.091.96 ± 0.09
vit huge 1.88±0.091.88$\pm$ 0.091.88 ± 0.09 1.92±0.071.92$\pm$ 0.071.92 ± 0.07
swin2 large 1.96±0.121.96$\pm$ 0.121.96 ± 0.12 1.98±0.111.98$\pm$ 0.111.98 ± 0.11
convnext large 2.10±0.062.10$\pm$ 0.062.10 ± 0.06 2.15±0.072.15$\pm$ 0.072.15 ± 0.07
vitg 14 1.76±0.081.76$\pm$ 0.081.76 ± 0.08 1.80±0.061.80$\pm$ 0.061.80 ± 0.06
resnet152 2.11±0.092.11$\pm$ 0.092.11 ± 0.09 2.09±0.112.09$\pm$ 0.112.09 ± 0.11
LLaVA(Zero-Shot) 2.95±0.332.95$\pm$ 0.332.95 ± 0.33 2.98±0.342.98$\pm$ 0.342.98 ± 0.34
LLaVA(fine-tuned) 1.70±0.221.70$\pm$ 0.221.70 ± 0.22 1.69±0.221.69$\pm$ 0.221.69 ± 0.22
@@
BLIP (fine-tuned) 1.94±0.051.94$\pm$ 0.051.94 ± 0.05 1.96±0.051.96$\pm$ 0.051.96 ± 0.05
BLIP (words fine-tuned) 1.92±0.061.92$\pm$ 0.061.92 ± 0.06 1.95±0.041.95$\pm$ 0.041.95 ± 0.04
GPT-4o(Zero-Shot) 2.612.612.61 2.632.632.63
Gemini-Flash 2.062.062.06 2.112.112.11
Humans 2.72±0.882.72$\pm$ 0.882.72 ± 0.88 2.71±0.862.71$\pm$ 0.862.71 ± 0.86
Chance(from distribution) 2.69±0.052.69$\pm$ 0.052.69 ± 0.05 2.50±0.062.50$\pm$ 0.062.50 ± 0.06
Chance(for zero-shot) 3.58±0.073.58$\pm$ 0.073.58 ± 0.07 3.33±0.073.33$\pm$ 0.073.33 ± 0.07
@@ 7.1 Training details
[…]
@@ 13 Image categories results
dinov2 large 59±359$\pm$ 359 ± 3 60±160$\pm$ 160 ± 1 47±147$\pm$ 147 ± 1 59±259$\pm$ 259 ± 2 51±251$\pm$ 251 ± 2
vit huge 64±364$\pm$ 364 ± 3 61±261$\pm$ 261 ± 2 48±248$\pm$ 248 ± 2 62±262$\pm$ 262 ± 2 52±252$\pm$ 252 ± 2
vit large 58±258$\pm$ 258 ± 2 58±258$\pm$ 258 ± 2 47±147$\pm$ 147 ± 1 59±259$\pm$ 259 ± 2 51±251$\pm$ 251 ± 2
swin2 large 61±261$\pm$ 261 ± 2 60±260$\pm$ 260 ± 2 47±147$\pm$ 147 ± 1 60±260$\pm$ 260 ± 2 52±252$\pm$ 252 ± 2
convnext large 57±257$\pm$ 257 ± 2 57±157$\pm$ 157 ± 1 46±146$\pm$ 146 ± 1 57±257$\pm$ 257 ± 2 50±050$\pm$ 050 ± 0
vitg 14 72±372$\pm$ 372 ± 3 70±370$\pm$ 370 ± 3 51±251$\pm$ 251 ± 2 68±368$\pm$ 368 ± 3 53±253$\pm$ 253 ± 2
resnet152 56±156$\pm$ 156 ± 1 55±255$\pm$ 255 ± 2 46±146$\pm$ 146 ± 1 56±256$\pm$ 256 ± 2 50±150$\pm$ 150 ± 1
LLaVA (Zero-Shot) 63±563$\pm$ 563 ± 5 66±466$\pm$ 466 ± 4 46±146$\pm$ 146 ± 1 52±252$\pm$ 252 ± 2 66±366$\pm$ 366 ± 3
LLaVA (fine-tuned) 72±272$\pm$ 272 ± 2 76±276$\pm$ 276 ± 2 54±254$\pm$ 254 ± 2 65±265$\pm$ 265 ± 2 69±369$\pm$ 369 ± 3
LLaVA (words fine-tuned) 79±279$\pm$ 279 ± 2 83±183$\pm$ 183 ± 1 54±154$\pm$ 154 ± 1 69±269$\pm$ 269 ± 2 73±173$\pm$ 173 ± 1
BLIP (fine-tuned) 59±159$\pm$ 159 ± 1 59±259$\pm$ 259 ± 2 48±148$\pm$ 148 ± 1 59±259$\pm$ 259 ± 2 52±252$\pm$ 252 ± 2
BLIP (words fine-tuned) 63±263$\pm$ 263 ± 2 66±266$\pm$ 266 ± 2 49±149$\pm$ 149 ± 1 61±261$\pm$ 261 ± 2 55±155$\pm$ 155 ± 1
GPT-4o (Zero-Shot) 757575 696969 505050 585858 767676
Gemini-Flash 737373 848484 535353 747474 828282

### v3

@@ 3.2 Assessing human performance
    Outlier Removal: Participants with >>10% of ratings having z-scores outside $\pm$1.96 were excluded
@@
  • Binary and Comparison Tasks: Excluded participants with >>3 inconsistent responses on repeated images
  • Range Task: Excluded participants if any single image received ratings differing by >>4 points, or if >>4 images showed differences exceeding 2 points
@@
  • Binary Task: The degree of self-consistency for repeated images was $84.2\pm 13.3\%$
@@
  • Comparison Task: The degree of self-consistency for repeated choices was $91.3\pm 14.8\%$
@@
Figure 3: Modifications rendered images less humorous. Each point compares the rating of image pairs (y-axis: original, x-axis: modified pair; total 1,271 pairs; line = identity). For the majority of images ($86.4\%$), the ratings for the original images were higher.
@@
Training $2,136$ 698 1,068 $0.79\pm 0.18$ $5.75\pm 1.42$ 1,068 $0.12\pm 0.16$ $3.62\pm 1.75$
Valid. $703$ 273 351 $0.78\pm 0.19$ $5.68\pm 1.40$ 351 $0.12\pm 0.17$ $3.65\pm 1.65$
testAll $706$ 300 352 $0.77\pm 0.19$ $5.60\pm 1.37$ 352 $0.14\pm 0.17$ $3.39\pm 1.68$
testOnlyPairs $600$ 300 300 $0.77\pm 0.19$ $5.62\pm 1.36$ 300 $0.13\pm 0.16$ $3.30\pm 1.68$
Total 3,542 1,271 1,771 $0.79\pm 0.20$ $5.70\pm 1.40$ 1,771 $0.13\pm 0.17$ $3.58\pm 1.72$
@@ 5 Results
Explanation Quality. Explanation Quality. Our human evaluation of VLM explanations revealed strong performance, particularly from larger models. On a 1–5 scale, Gemini’s explanations were rated highest for both identifying the source of humor ($4.1\pm 0.6$) and for overall accuracy ($4.24\pm 0.47$). GPT-4o also performed well (Why Funny: $3.9\pm 0.7$, Accuracy: $4.12\pm 0.63$), followed by LLaVA (Why Funny: $3.5\pm 0.9$, Accuracy: $3.57\pm 0.95$). These results suggest a robust, though not perfect, ability to articulate visual humor beyond simple keyword matching. Automated scores are detailed in Appendix 8.
@@
Model Name testAllSet RMSE $\pm SD$ testOnlyPairs RMSE $\pm SD$
dinov2 large $1.98\pm 0.08$ $1.96\pm 0.09$
vit huge $1.88\pm 0.09$ $1.92\pm 0.07$
swin2 large $1.96\pm 0.12$ $1.98\pm 0.11$
convnext large $2.10\pm 0.06$ $2.15\pm 0.07$
vitg 14 $1.76\pm 0.08$ $1.80\pm 0.06$
resnet152 $2.11\pm 0.09$ $2.09\pm 0.11$
LLaVA(Zero-Shot) $2.95\pm 0.33$ $2.98\pm 0.34$
LLaVA(fine-tuned) $1.70\pm 0.22$ $1.69\pm 0.22$
@@
BLIP (fine-tuned) $1.94\pm 0.05$ $1.96\pm 0.05$
BLIP (words fine-tuned) $1.92\pm 0.06$ $1.95\pm 0.04$
GPT-4o(Zero-Shot) $2.61$ $2.63$
Gemini-Flash $2.06$ $2.11$
Humans $2.72\pm 0.88$ $2.71\pm 0.86$
Chance(from distribution) $2.69\pm 0.05$ $2.50\pm 0.06$
Chance(for zero-shot) $58\pm 0.07 3.3$$33\pm 0.07$
@@ 7.1 Training details
Models were trained using the Adam optimization algorithm with weight decay. A hyperparameter grid search was conducted across learning rates in the set $\{0.01,0.001,0.0001,0.00001\}$, batch sizes in the set $\{4,8,16\}$, and weight decay parameters in the set $\{0.1,0.01,0.001\}$. Model training proceeded for a fixed number of 10 epochs, with periodic checkpoints. For the final evaluation on the unseen test set, we used the model iteration exhibiting optimal performance on the validation cohort. We used cross-entropy loss for the Binary classification and Comparison tasks while mean square loss was used for the Regression task. For all architectures except GPT-4o and Gemini-Flash, we fine-tuned the models(Pretraining details are mentioned in Section 4.1). For LlaVA we did lora fine-tuning instead of full fine-tuning. To ensure statistical robustness, each experiment was conducted 5 times for all the experiments except for GPT-4o and Gemini-Flash which were run only one time.
@@ 13 Image categories results
dinov2 large $59\pm 3$ $60\pm 1$ $47\pm 1$ $59\pm 2$ $51\pm 2$
vit huge $64\pm 3$ $61\pm 2$ $48\pm 2$ $62\pm 2$ $52\pm 2$
vit large $58\pm 2$ $58\pm 2$ $47\pm 1$ $59\pm 2$ $51\pm 2$
swin2 large $61\pm 2$ $60\pm 2$ $47\pm 1$ $60\pm 2$ $52\pm 2$
convnext large $57\pm 2$ $57\pm 1$ $46\pm 1$ $57\pm 2$ $50\pm 0$
vitg 14 $72\pm 3$ $70\pm 3$ $51\pm 2$ $68\pm 3$ $53\pm 2$
resnet152 $6\pm 1 55$$5\pm 2$ $46\pm 1$ $56\pm 2$ $50\pm 1$
LLaVA (Zero-Shot) $3\pm 5 66$$6\pm 4$ $46\pm 1$ $52\pm 2$ $66\pm 3$
LLaVA (fine-tuned) $72\pm 2$ $76\pm 2$ $54\pm 2$ $65\pm 2$ $69\pm 3$
LLaVA (words fine-tuned) $79\pm 2$ $83\pm 1$ $54\pm 1$ $69\pm 2$ $73\pm 1$
BLIP (fine-tuned) $59\pm 1$ $59\pm 2$ $48\pm 1$ $59\pm 2$ $52\pm 2$
BLIP (words fine-tuned) $3\pm 2 66$$6\pm 2$ $49\pm 1$ $61\pm 2$ $55\pm 1$
GPT-4o (Zero-Shot) $75$ $69$ $50$ $58$ $76$
Gemini-Flash $73$ $84$ $53$ $74$ $82$